OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through point (1, -2) and has a slope of 1/3

OpenStudy (anonymous):

the equation of a line is:y-y1=m(x-x1) where m is the slope. y-(-2))=1(x-1)/3 y+2=(x-1)/3 3(y+2)=x-1 3y+6=x-1 3y+5=x x-3y=5 which is the equation of the line.

OpenStudy (anonymous):

We can write : \[y=\frac13x+b\] To find b we substitute with the point (1,-2) like this : \[-2=\frac13\times 1 +b\] So : \[b=-2-\frac13=-\frac73\] Therefore : \[y=\frac13x-\frac73\]

OpenStudy (anonymous):

oh i did a mistake actually it will be 3y+6=x-1 3y+7=x so x-3y=7 is the equation..